OpenStudy (anonymous):

A Math test has a mean of 39 and standard deviation of 18.0. Find the corresponding z scores for: I. a test score of 31 II. a test score of 100 Please show all of your work

OpenStudy (sirm3d):

the z score formula is \[z=\frac{ obsv - mean }{ std.dev. }\] i'll leave the computation to you. post your answers here.

OpenStudy (anonymous):

what is obsv?

OpenStudy (anonymous):

z=31-39/18.0 correct?

OpenStudy (anonymous):

z=100-39/18.0 = 3.4

OpenStudy (anonymous):

z=31-39/18.0=.44

OpenStudy (anonymous):

is this right so far?

OpenStudy (sirm3d):

there is a negative sign

OpenStudy (sirm3d):

I. z = -0.44 II. z = 3.4

OpenStudy (anonymous):

okay so I was on the right track then just forgot the negative sign

OpenStudy (sirm3d):

just be careful with the signs.

OpenStudy (anonymous):

Thank you so much for the help.